if $1000 is equal to 184,000 naira. how much $ is equal to 100,000 naira?
Answer by Theo(13342)

your equation is:

1000 / 184,000 = x / 100,000
multiply both sides of this equation by 100,000 to get:
1000 * 100,000 / 184,000 = x
solve for x to get:
x = 543.4782609 dollars.

another way to look at it is:

1000 dollars is equal to 184,000 naira.

divide both sides of that equation by 184,000 and you get:

1000 / 184,000 dollars is equal to 184,000 / 184,000 naira.

simplify to get:

.0054347826 dollars equals 1 naira.

100,000 naira would therefore be equal to 100,000 * .0054347826 which makes 100,000 naira equal to 543.347826 naira.

543.347826 is the same number as 543.34782609, with the difference being rounding caused by the limitation of the number of digits that can be displayed on the calculator.